OpenStudy (austinl):

What is the next number in this series? 6, 14, 36, 98, 276, ?

OpenStudy (austinl):

No worries by the way, I have been posting riddles for the fun of it this morning. It is just for our own fun.

OpenStudy (austinl):

@terenzreignz

terenzreignz (terenzreignz):

no idea still :3

OpenStudy (austinl):

This one is a hard one!

OpenStudy (rane):

is the next no' 794

OpenStudy (rane):

@austinL

OpenStudy (austinl):

VERY GOOD!!!!!

OpenStudy (rane):

r u serious ?

OpenStudy (austinl):

YEP!!

OpenStudy (rane):

yeah!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (austinl):

794 The nth term in the series is given by the sum of x^n for x in the range 1 to 3, i.e. 1^n + 2^n + 3^n. Thus the first term is 1^1 + 2^1 + 3^1 = 1 + 2 + 3 = 6, the second term is 1^2 + 2^2 + 3^2 = 1 + 4 + 9 = 14, and so on. The sixth term is then 1^6 + 2^6 + 3^6 = 1 + 64 + 729 = 794.

OpenStudy (rane):

lol that's wht I did

OpenStudy (rane):

yes!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (austinl):

Good job, this was a very hard one!

OpenStudy (rane):

thanks :)

OpenStudy (rane):

there is another way @austinL

OpenStudy (austinl):

is there?

OpenStudy (rane):

yup

OpenStudy (rane):

for the ist term : (6*3) = 18-4 = 14 2nd term : (14*3) = 42 - 6 = 36 3rd term : (36*3) = 108 - 10 = 98 4th term : (98*3) = 294 - 18 = 276 5th term : (276*3) = 828 - 34 = 794 the pattern is that u have to multiply each term by 3. nd times the ist term by 4. now: 4 + 2 =6, 6+2+2 = 10, 10+2+2+2+2 = 18 (u have double the amount of 2's every time)

OpenStudy (austinl):

Wow, very cool!

OpenStudy (rane):

YUP I KNOW